OpenStudy (anonymous):

right triangular prism with base legs 8 cm and 10 cm and height 20 cm and it was volume

OpenStudy (anonymous):

volume right triangular prism=1/2*length*width*height

OpenStudy (anonymous):

ok so it was 800cm

OpenStudy (anonymous):

am i right or not?

OpenStudy (anonymous):

right

jhonyy9 (jhonyy9):

volum = area of base time height

OpenStudy (anonymous):

it was 400? or 800

jhonyy9 (jhonyy9):

- so area of a right triangle equal ?

OpenStudy (anonymous):

400

OpenStudy (anonymous):

which one is correct 800 or 400 some peolpe said it was 800 and some people said it was 400

jhonyy9 (jhonyy9):

can you make the calcules ?

jhonyy9 (jhonyy9):

- so how many will be area of base ?

OpenStudy (anonymous):

base legs is 8

jhonyy9 (jhonyy9):

yes ... and ?

OpenStudy (anonymous):

and 10

jhonyy9 (jhonyy9):

yes and area of base will be , how many ?

OpenStudy (anonymous):

it was 400

jhonyy9 (jhonyy9):

no

OpenStudy (anonymous):

800

jhonyy9 (jhonyy9):

i like to learn you how you need to do these calcules

OpenStudy (anonymous):

800

jhonyy9 (jhonyy9):

first you need to calcule the area of base

OpenStudy (anonymous):

the area base is 80

jhonyy9 (jhonyy9):

what in a right triangle is base time height divide by 2

jhonyy9 (jhonyy9):

so than area of base equal 80/2 =40

jhonyy9 (jhonyy9):

so than the volum will be 40 time the height of prisme what is 20

jhonyy9 (jhonyy9):

so than

jhonyy9 (jhonyy9):

so than

jhonyy9 (jhonyy9):

40*20=800

jhonyy9 (jhonyy9):

than 800 is the right answer

OpenStudy (anonymous):

thanks

jhonyy9 (jhonyy9):

ok ?

jhonyy9 (jhonyy9):

you have understood it sure ?

OpenStudy (anonymous):

yeah i understood now :) thanks so much

jhonyy9 (jhonyy9):

ok than can you say me how many will be the volum if there are a piramid and not prisme ?

jhonyy9 (jhonyy9):

- with this dimensions

OpenStudy (anonymous):

what do you mean?

jhonyy9 (jhonyy9):

- so if you will say me i will believe you that you know how you ned to do these calcules

jhonyy9 (jhonyy9):

- i have thought that there is a piramide and not a prisme

jhonyy9 (jhonyy9):

- so than what will be the volum for this piramide ?

OpenStudy (anonymous):

it was 800

jhonyy9 (jhonyy9):

- no this is a new exercise for you

OpenStudy (anonymous):

then what is question

jhonyy9 (jhonyy9):

- for your first exercise the right answere is 800 yes

jhonyy9 (jhonyy9):

but do you know the formula for one piramide ?

jhonyy9 (jhonyy9):

formule for volum of a piramide

jhonyy9 (jhonyy9):

do you like to learn it ?

jhonyy9 (jhonyy9):

ok bye good luck